OpenStudy (anonymous):

This exercise deals with logarithmic scale. If one earthquake is 20 times as intense as another, how much larger is its magnitude on the Richter scale?

OpenStudy (anonymous):

Not familiar, some formula stuff here http://en.wikipedia.org/wiki/Moment_magnitude_scale

OpenStudy (anonymous):

we need richter scale formula for this one

OpenStudy (anonymous):

let me find it

OpenStudy (anonymous):

actually we probably don't. it is logarithmic scale, so answer is \[\log(20)\]

OpenStudy (anonymous):

It's on the link I gave, I thought...

OpenStudy (anonymous):

i think the answer is about 3

OpenStudy (anonymous):

i think its supposed to be 1.3

OpenStudy (anonymous):

m=log I/s

OpenStudy (anonymous):

where I is intensity and s is standard earthquake

OpenStudy (anonymous):

m is magnitude

OpenStudy (anonymous):

if intensity is of the first one is I then the richter scale is something like \[\log(I)\] then 20 times as intense gives \[\log(20I)=\log(20)+\log(I)\] and \[\log(20)=2.99...\] so i think it is about 3 more

OpenStudy (anonymous):

are you sure of the answer 1.3 because i do not get that. the dividing by s part is not important. it is just a logarithmic scale

OpenStudy (anonymous):

ya the professor says we should get log 20 = 1.3

OpenStudy (anonymous):

oooooooooooooooooh because si am a moron

OpenStudy (anonymous):

i was using the natural log and not the common log.

OpenStudy (anonymous):

it is \[\log(20)\] as i said, but \[\log(20)=1.3...\]

OpenStudy (anonymous):

ah i see

OpenStudy (anonymous):

that is correct. i had the wrong log in my calculator.

OpenStudy (anonymous):

sorry for the confustion

OpenStudy (anonymous):

thanks for clearing that up